OpenStudy (anonymous):

Express the complex number in trigonometric form. -3 + 3 square root of 3i

OpenStudy (anonymous):

\[put r \cos theta=-3, r \sin \theta=3\sqrt{3}\] square and add, \[r ^{2}\left( \cos ^{2}\theta+\sin ^{2}\theta\right)=9+27=36\] \[r ^{2}=36,r=6,\] dividing, \[\tan \theta=\frac{ 3\sqrt{3} }{ -3}=-\sqrt{3}=-\tan \frac{ \pi }{ 3 }=\tan \left( \pi-\frac{ \pi }{3 } \right)\] \[\tan \theta=\tan \frac{ 2\pi }{ 3 }\] \[\theta=\frac{ 2\pi }{3 }\] \[-3+3\sqrt{3}\iota=6\left( \cos \frac{ 2\pi }{ 3 }+\iota \sin \frac{ 2\pi }{3 } \right)\]